Ecker Company reports $1,800,000 of net income and declares $252,000 of cash dividends on its preferred stock for the year. At year-end, the company had 260,000 weighted-average shares of common stock.

1. What amount of net income is available to common stockholders?
2. What is the company’s basic EPS?
Answer #1

1.Net income available to common stockholders=Net income-Cash dividends on preferred shares

=(1,800,000-252,000)=$1,548,000

2.Basic EPS=(Net income-Cash dividends on preferred shares)/weighted-average shares of common stock

=(1,800,000-252,000)/260,000

=$5.95 per share(Approx).